OpenStudy (anonymous):

Use the following equation to answer the question below: Fe2O3 + 3CO(g) -> 2Fe(s) + 3CO2(g) What volume of CO2 gas would be produced at 25C and 745 mmHg if 125 g of Fe2O3 reacted?

OpenStudy (matt101):

We already know we're going to need to use the ideal gas law, PV=nRT, at some point because we need to go from moles to volume. The first thing to do is find moles of CO2, and we know from the equation that there are 3x as many moles of CO2 produced as there are moles of Fe2O3 consumed. We can find moles of Fe2CO3 because we know the mass - just divide 125 g by the molar mass of Fe2CO3 (160 g/mol) to get 0.781 mol. There is 3x as much CO2, or 2.343 mol. Now we can plug all our values into PV=nRT to solve for volume. Remember to use the right units! P = 745 mmHg = 0.980 atm n = 2.343 mol R = 0.08206 L・atm / (K・mol) T = 25 C = 298 K V = nRT/P = (2.343)(0.08206)(298)/0.98 = 58.5 L Please ask if you have any questions!